In a certain population, 16% of the people have Rh-negative blood. A blood bank serving this population receives 95 blood donors on a particular day. (Use the normal approximation. Round your answers to four decimal places.)

a) What is the probability that 10 or fewer are Rh-negative?

b)What is the probability that 15 to 20 (inclusive) of the donors are Rh-negative?

c)What is the probability that more than 80 of the donors are Rh-positive?
Answer #1

Proportion of people have Rh- negative blood P=0.16 n = 95 np - 95(0.16) -15-2 (1:9=1-P) = 79.8 na z 95(1-0-16) greater than95 x0.16X 0-84 Os3-573 x Norm (1,6) Random Variable that represents number of people has Rh- negative. @ P(xcio convert = int© PC15 CX620) = 15-15-2 ? ( 15 L Z L 20-15.2 3.573 (3.573) P(-0.06 LZL 1,34) =P(z 21.34 -P(Z2-0.06) P(Z <1,34) left tail areaP (more than so of the donors are Rh positive) have As we know that 16% of the people Rh- negative blood of the people have tx~ Random Variable that represents of people have Rh positive number P(x>807 Convert into normal random Variable Z score = P(